OpenStudy (anonymous):

The probability that is a fine day is 3/7. If it is a fine day Joe will cycle to work 70% of the time and drive to work 20% of the time. He can also take the train to work. If it is a wet day, Joe will cycle to work 1/9 of the time and drive to work 5/9 time. Joe can also take the train. If Joe works 315 days of the year, find how many days he is likely to drive to work.

OpenStudy (kc_kennylau):



OpenStudy (anonymous):

I am completing the above work \[ \frac{3}{7} \, \frac 1 5+\frac47 \, \frac 5 9=\frac{127}{315}=0.403 \]

OpenStudy (anonymous):

About 40% of the time

OpenStudy (anonymous):

127 days